# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

### 05 Apr 2024 v2

@@ 5.5. Subset Selection Algorithms
(15) $\mathop{\mathrm{min}}\limits_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|.$
@@
where $\hat{\theta}^{T}=\arg\min_{\theta}{\sum_{(q,y)\in S}{l(y,f(q,\theta))}}$ is the final proficiency estimate when the test ends $(t=T)$ with the corresponding $T$ responses. In contrast to previous sequential selection methods, it essentially does not require perfect selection at each step, but rather emphasizes the accuracy of the final estimate. From a global perspective, CAT essentially is a Subset Selection problem (Miller, 2002), a fundamental challenge in machine learning and optimization. It revolves around choosing a subset of elements $S$ from a larger set $\mathcal{Q}$ that optimizes a particular objective function $F(S)$ while adhering to specific constraints. However, we cannot directly solve the above optimization problem due to the following main challenge: The true proficiency of the examinee, denoted by $\theta_{0}$, is unknown. It is not available in the dataset, which prevents us from directly optimizing or designing the question selection algorithm. To address this, some researchers have tried to approximate the original objective. Mujtaba and Mahapatra (2021) use the standard error of measurement (SEM) to approach the optimization objective. The SEM provides a measure of confidence in an estimate from a test. They propose using SEM as a new target to measure the accuracy of proficiency estimation: $\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|\Rightarrow\min_{|S|=T}\sqrt{\frac{1}{\sum_{i\in S}{\mathcal{I}_{i}(\hat{\theta}^{T})}}},$ where $\mathcal{I}_{i}(\hat{\theta})$ is the information (e.g., Fisher Information) for a question $i$ for an examinee with proficiency $\hat{\theta}$, thus the objective function $F(S)=\sqrt{\frac{1}{\sum_{i\in S}{\mathcal{I}_{i}(\hat{\theta}^{T})}}}$. At each step, it uses multi-objective evolutionary algorithms to obtain the set of Pareto-optimal solutions (Deb, 2011) by maximizing precision and minimizing the number of questions.
Figure 8. Illustration of the subset optimization problem, adapted from (Zhuang et al., 2023): Selecting subset $S$ to cover the bank $Q$. Rectangles represent different questions, with $w(i,j)$ measuring the similarity of question pair.
@@
To achieve a more general and scalable framework, Zhuang et al. (2023) propose a bounded estimation CAT framework (BECAT), which redefines solve the selection problem in a data summary manner. Due to the unknown $\theta_{0}$, they first find its theoretical approximation as the new target for optimization: The examinee’s one proficiency estimate $\theta^{*}$, estimated by their full responses to the entire question bank $\mathcal{Q}$, is a theoretical approximation of the true proficiency, i.e., $\theta^{*}\approx\theta_{0}$. This finding allows the selection algorithm to aim to approach $\theta^{*}$ instead of the unknown $\theta_{0}$: Select a subset of questions $S\subseteq\mathcal{Q}$, so that the examinee’s proficiency is estimated only on the subset $S$ while still (approximately) converging to the $\theta^{*}$ (i.e., the estimate that would be obtained if optimizing on the full responses to $\mathcal{Q}$).
$\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|\Rightarrow\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta^{*}\|$ $\Rightarrow\min_{|S|=T}\max_{\theta\in\Theta}\left\|\sum_{(q,y)\in S}{\gamma\nabla l(y,f(q,\theta))}-\sum_{(q,y)\in\mathcal{Q}}{\nabla l(y,f(q,\theta))}\right\|$
$\Rightarrow\mathop{\mathrm{min}}\limits_{|S|=T}\mathop{\mathrm{max}}\limits_{\theta\in\Theta}\sum_{i\in\mathcal{Q}}\mathrm{min}_{j\in S}\|\nabla l_{i}(\theta)-\nabla l_{j}(\theta)\|$
(16) $\Rightarrow\mathop{\mathrm{max}}\limits_{|S|=T}\sum_{i\in\mathcal{Q}}\mathrm{max}_{j\in S}\;w(i,j),$
@@
where $w(i,j)\triangleq d-\mathrm{max}_{\theta\in\Theta}{\|\nabla l_{i}(\theta)-\nabla l_{j}(\theta)\|}$ is the gradient similarity between question pair $(q_{i},q_{j})$ for this examinee, thus the objective function $F(S)=\sum_{i\in\mathcal{Q}}\mathrm{max}_{j\in S}\;w(i,j)$. The core of BECAT’s subset selection strategy is to find a subset $S$ of size $T$ that maximizes the coverage of $\mathcal{Q}$, quantified by the similarity measure $w(i,j)$ between each question $i\in\mathcal{Q}$ and its closest counterpart $j\in S$. This approach, represented in Figure 8, essentially seeks the most representative questions, aligning with prior selection algorithms but under a new, more rigorous theoretical framework. Given the NP-Hard nature of this optimization, BECAT employs a submodular function approximation. A simple greedy algorithm can finally solve this subset selection problem, with BECAT ensuring that the estimate error remains upper-bounded at each step. The subset selection problem in CAT is a fresh direction with significant potential. This method offers a universal and global framework for question selection, applicable across various complex CDMs that can utilize gradient-based estimations, including neural network models. Crucially, it also provides a theoretical basis for estimation error analysis.
Table 1. Comparison of Different Question Selection Algorithms in CAT
@@
Statistical    Algorithms
Simple implementation and efficient operation
@@
Dependent on IRTs and requires expert knowledge for design
Active Learning
Model-agnostic and flexible
@@
Neglect the nuanced information within CDM parameters
Reinforcement Learning
Automatic generation of selection algorithm; Sequential Decision Making
@@
Incurs additional training costs and potential bias from data-driven selection
Meta Learning
Automatic generation of selection algorithm; Fast Adaptation
@@
Incurs additional training costs and potential bias from data-driven selection
Subset Selection
@@
Strong theoretical guarantees for estimation accuracy
Faces challenges in the initial stages of CAT
Discussion 5.
@@
Table 1 compares these five categories of selection algorithms in CAT systems, highlighting the generality and interpretability of each category, along with their main advantages and limitations. This overview assists researchers in identifying the most suitable algorithm for their CAT applications, balancing efficiency and complexity (a summary of the performance of different methods is presented in Table 3). It is noteworthy that, despite the superior performance demonstrated by the latest machine learning and deep learning methods, they do not currently supplant traditional statistical approaches. Particularly in testing scenarios that prioritize interpretability or efficiency, statistical methods remain predominant.
6. Question Bank Construction
@@
To develop a high-quality CAT, the foundational step is to construct a high-quality question bank with a substantial number of questions. The construction of a question bank can be decomposed into two main stages as shown in Figure 9: Question Characteristics Analysis and Question Bank Development: (1) Question Characteristics Analysis first detailedly examines the properties and attributes of potential questions. Then, (2) Question Bank Development assembles the final question bank $\mathcal{Q}$ from the analyzed questions.
6.1. Question Characteristics Analysis
@@
The first stage, question characteristics analysis, involves a detailed examination of the properties and attributes of potential questions, e.g., difficulty, discrimination, and the knowledge concepts required to answer the question (i.e., Q-matrix). To illustrate, when selecting questions based on Fisher Information, one must leverage pre-calibrated parameters like difficulty ($\beta_{j}$), discrimination ($\alpha_{j}$), and guessing factor ($c_{j}$), alongside the current proficiency estimate, to compute the Information value $\mathcal{I}_{j}(\theta)$ for each question $j$. The methods of characteristics analysis can be categorized into three main approaches: expert-based, statistic-based, and deep learning-based methods.
6.1.1. Expert-based Characteristics Annotation
In expert-based annotation, domain experts assess question parameters, a practice seen in online CAT systems like SIETTE (Conejo et al., 2004) and GenTAI (López-Cuadrado et al., 2006). Effective expert estimation often involves structured questionnaires (López-Cuadrado et al., 2008) that clearly outline the task and instructions. Follow-up discussions may refine these assessments, especially when expert opinions diverge. Results are aggregated, using averages for continuous attributes or voting for discrete ones, and sometimes applying more sophisticated statistical methods like the Consensus Theory algorithm (Kozierkiewicz-Hetmańska and Poniatowski, 2014) for expert consensus. While this approach is reliable for estimating difficulty, it can be less so for other parameters like discrimination and guessing factor (Lopez-Cuadrado et al., 2009). Furthermore, expert judgments can sometimes be subjective, potentially leading to inaccuracies in annotations, particularly when expert numbers are limited or their views are inconsistent.
6.1.2. Statistic-based Characteristics Annotation
@@
The statistic-based method for annotating question characteristics requires gathering responses from a large group of examinees. Despite its resource-intensive nature, involving pre-testing with examinees, it utilizes parameter estimation methods or gradient descent on CDM (e.g., IRT) to estimate question parameters (De Champlain, 2010). In some psychological theories, like Classic Test Theory (CTT), question difficulty is calculated as the proportion of correct responses within examinees (Magno, 2009; DeVellis, 2006), while discrimination is derived from performance disparities between higher and lower ability examinees (Chang and Yang, 2009). The Q-matrix is another crucial characteristic of questions. It is a binary matrix that indicates which knowledge concepts are associated with a question. Numerous researchers have attempted to employ some data-driven approaches to learn the Q-matrix (Liu et al., 2013; Chen et al., 2015b). Data-driven methods like Boolean Matrix Factorization (BMF) (Sun et al., 2014) have been applied to infer it from examinee responses. The examinee response matrix ${R}$ is approximated by the element-wise multiplication of the logical complements of the examinee proficiency matrix ${A}$ and the transpose of the Q-matrix ${Q}$:
(17) ${R}=\overline{\overline{{A}}\odot{Q}^{T}}.$
The Q-matrix can then be learned by solving this equation, given the examinee response matrix and an initial estimate of the examinee proficiency matrix. An alternate recursive algorithm with a Heaviside step function is used to address the NP-hard nature of BMF and the non-invertibility of Boolean operations. Further enhancements to this method include additional judgment and processing steps to ensure convergence and robustness, as well as refinement procedures to avoid local optima in Q-matrix estimation (Xiong et al., 2022).
Figure 9. For high-quality question bank construction, two main stages are needed: question characteristics analysis and question bank development.
6.1.3. Deep Learning-based Characteristics Annotation
@@
With the rise of Natural Language Processing (NLP), there has been an increasing trend in recent years to directly use the textual information of questions to analyze various attributes. For difficulty prediction, attention-based CNN models and domain adaptation strategies have been used to evaluate reading questions and medical question complexity (Huang et al., 2017; Qiu et al., 2019; Huang et al., 2021). For knowledge concept (Q-matrix) prediction, which typically exhibits a hierarchical structure, a Hierarchical attention-based Recurrent Neural Network has been proposed (Huang et al., 2019, 2022). Lei et al. (2021) further take into account the multi-modal features of questions, such as images and formulas. Pre-trained NLP models have also proven effective for automated question analysis (Yin et al., 2019; Ning et al., 2023).
6.2. Question Bank Development
@@
The second stage, question bank development, involves the actual assembly of the question bank from the analyzed questions of the first stage. This process should aim to create a balanced and varied bank that can cater to different levels of proficiency and different areas of knowledge (Revuelta and Ponsoda, 1998; Reckase, 2010; Segall, 2005). This stage is critical as the quality of the question bank directly impacts the CAT system’s effectiveness. According to different usage scenarios, the approaches to developing a question bank can be categorized into the following three aspects.
6.2.1. Question Bank Blueprint Design
@@
The goal of designing a blueprint is to generate an optimal framework for a question bank. This blueprint outlines the distribution of questions based on various statistical and non-statistical attributes associated with the questions. Reckase (2010) initially conduct a theoretical analysis of the characteristics of an optimal question bank. In a CAT system using a 1PL-IRT model with a maximum Fisher information selection algorithm, an optimal question bank is characterized by questions whose difficulty level exactly matches the estimated proficiency for each selection. However, achieving an exact match is often unfeasible. Thus, they propose a method known as the bin-and-union, which allows for a maximum deviation $r$ between optimal difficulty and estimated proficiency, is used to derive the question bank size and $r$-optimal distribution. Further research has extended these methods for large-scale CAT systems (He and Reckase, 2014) and continuous new question pretesting (Van Der Linden et al., 2000).
6.2.2. Question Bank Assembly
@@
While the design of a question bank blueprint assumes the absence of actual questions and focuses on creating an optimal framework, the bank assembly is a process that involves generating question banks from an existing master bank according to specific requirements. Way et al. (2005) delve into the issues related to the development and maintenance of a master bank, including the constraints necessary to ensure the assembled question bank meets desired specifications. van der Linden et al. (2006) propose a mixed-integer programming model to create a question bank that satisfies all content specifications and maximizes information at selected ability values. This method, illustrated using the Law School Admission Test’s master bank, focuses on distributing estimates for the examinee population. Belov and Armstrong (2009) further introduces two significant problems related to question bank design: the direct problem and the inverse problem. The direct problem involves creating a bank that meets specific measurement requirements, such as content coverage and measurement precision, while minimizing question exposure. Conversely, the inverse problem involves determining the minimum number of questions needed to achieve the above specific requirements. A Monte Carlo method is introduced to balance measurement accuracy and question exposure.
6.2.3. Question Bank Rotating
@@
Rotating the question bank refers to dividing a master bank into a set of smaller question banks, which may contain overlapping elements. This technique allows for the rotation of these smaller question banks among examinees, thereby ensuring a balanced exposure rate of test questions (Stocking and Swanson, 1998; Ariel et al., 2004; Swanson and Stocking, 1993). Ariel et al. (2004) proposes a method of dividing a master bank into smaller banks based on Gulliksen’s matched random subtests method (Gulliksen, 2013) to prevent issues of over- or underexposure in adaptive testing. The Weighted Deviation Model (WDM) (Swanson and Stocking, 1993) is employed to manage the degree of overlap between these banks, maintaining their representativeness while preventing question overexposure. This balance is critical to preserve the integrity of the rotation system and ensure consistent measurement across different test versions.
Discussion 6.
@@
Think of the entire bank development process as creating and managing a library. The blueprint design (Section 6.2.1) is like the architectural plan for the library, defining where each section (e.g., fiction, non-fiction) will be located. The assembly process (Section 6.2.2) resembles acquiring books from suppliers based on specific demands. Rotating the question bank (Section 6.2.3) is similar to periodically rotating the books on display. Even though the library has a vast collection, only a subset is displayed prominently at any given time. This rotation ensures that different books get exposure, and library visitors encounter a variety of books over time. Similarly, rotating the question bank in CAT ensures a balanced exposure of test questions, enhancing the fairness and robustness of the testing process.
The construction of a high-quality question bank introduced in this section is not just a prerequisite for CAT, but also a continuous process. It requires regular updates and refinements to ensure the relevance and effectiveness of the adaptive testing system.
7. Test Control of CAT

### 09 Mar 2026 v3

@@ 5.5 Subset Selection Algorithms
$\mathop{\mathrm{min}}\limits_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|,$ (16)
where $\hat{\theta}^{T}=\arg\min_{\theta}{\sum_{(q,y)\in S}{\ell(y,f(q,\theta))}}$ is the final proficiency estimate when the test ends with the corresponding $T$ responses. In contrast to previous sequential selection methods, it essentially doesn’t require perfect selection at each step, but rather emphasizes the accuracy of the final estimate.
@@
From a global perspective, CAT essentially is a Subset Selection problem [114], a fundamental challenge in machine learning and optimization. It revolves around choosing a subset of elements $S$ from a larger set $\mathcal{Q}$ that optimizes a particular objective function $F(S)$ while adhering to specific constraints. However, we cannot directly solve the above optimization problem due to the following main challenge: The true proficiency of the examinee, denoted by $\theta_{0}$, is unknown. It is not available in the dataset, which prevents us from directly optimizing or designing the question selection algorithm. To address this issue, some researchers have developed heuristic methods. For example, Mujtaba et al. [119] use the standard error of measurement as the objective $F(S)$, which provides a measure of confidence in an estimate from a test. At each step, it uses multi-objective evolutionary algorithms to obtain the set of Pareto-optimal solutions [42] by maximizing precision and minimizing the number of questions. Recently, for AI model evaluation, clustering techniques (e.g., K-means) have been used to select representative subsets $S$ from benchmarks [125].
@@
To develop a more general and scalable CAT framework, Zhuang et al. [191] propose BECAT, which reformulates the question selection problem in a data summary manner. Since the true proficiency $\theta_{0}$ is unobservable, they approximate it using $\theta^{*}$: the proficiency estimated from an examinee’s full responses to the entire question bank $\mathcal{Q}$, i.e., $\theta^{*}\approx\theta_{0}$. This approximation enables the selection algorithm to target $\theta^{*}$ instead of the unknown $\theta_{0}$: Select a subset of questions $S\subseteq\mathcal{Q}$ such that the estimated proficiency based on $S$ closely approximates $\theta^{*}$ (i.e., the estimate that would be obtained if optimizing on the full responses to $\mathcal{Q}$).
$\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|\Rightarrow\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta^{*}\|$
$\Rightarrow$ $\min_{|S|=T}\max_{\theta\in\Theta}\Big\|\sum_{(q,y)\in S}{\gamma\nabla\ell(y,f(q,\theta))}-\sum_{(q,y)\in\mathcal{Q}}{\nabla\ell(y,f(q,\theta))}\Big\|$
$\Rightarrow$ $\mathop{\mathrm{min}}\limits_{|S|=T}\mathop{\mathrm{max}}\limits_{\theta\in\Theta}\sum_{i\in\mathcal{Q}}\mathrm{min}_{j\in S}\|\nabla\ell_{i}(\theta)-\nabla\ell_{j}(\theta)\|$
$\Rightarrow$ $\mathop{\mathrm{max}}\limits_{|S|=T}\sum_{i\in\mathcal{Q}}\mathrm{max}_{j\in S}\;w(i,j),$ (17)
@@
where $w(i,j)\triangleq d-\mathrm{max}_{\theta\in\Theta}{\|\nabla\ell_{i}(\theta)-\nabla\ell_{j}(\theta)\|}$ is the gradient similarity between question pair $(q_{i},q_{j})$ for this examinee, thus the objective function $F(S)=\sum_{i\in\mathcal{Q}}\mathrm{max}_{j\in S}\;w(i,j)$. The core of BECAT’s subset selection algorithm is to find a subset $S$ of size $T$ that maximizes the coverage of $\mathcal{Q}$, quantified by the similarity measure $w(i,j)$. This approach (Figure 6) essentially seeks the most representative questions, aligning with prior selection algorithms but under a new, more rigorous theoretical framework.
Given the NP-Hard nature of this optimization, BECAT employs a submodular function approximation. A simple greedy algorithm can finally solve this subset selection problem, with BECAT ensuring that the estimate error remains upper-bounded at each step. The subset selection problem in CAT is a fresh direction with significant potential. This method offers a universal framework for question selection, applicable across various complex measurement models that can utilize gradient-based estimations, including neural network models.
TABLE I: Comparison of Different Question Selection Algorithms in CAT
@@
Statistical    Algorithms Simple implementation and efficient operation Dependent on IRTs and requires expert knowledge for design
@@
Active Learning Model-agnostic and flexible Neglect the nuanced information within measurement model parameters
@@
Reinforcement Learning Automatic generation of selection algorithm; Sequential Decision Making Incurs additional training costs and potential bias from data-driven selection
@@
Meta Learning Automatic generation of selection algorithm; Fast Adaptation Incurs additional training costs and potential bias from data-driven selection
@@
Subset Selection Strong theoretical guarantees for estimation accuracy Faces challenges in the initial stages of CAT
@@
Discussion: It is noteworthy that, despite the superior performance demonstrated by the latest machine learning and deep learning approaches [181], they have not yet replaced traditional statistical approaches in practice. Particularly in testing scenarios that prioritize interpretability or efficiency, statistical methods remain predominant. In the Appendix, we compare these five categories of selection algorithms in CAT systems, highlighting the generality and interpretability of each category, along with their main advantages and limitations. This overview assists researchers in identifying the most suitable algorithm for their CAT applications, balancing efficiency and complexity.
6 Question Bank Construction
@@
To develop a high-quality CAT, the foundational step is to construct a high-quality question bank. The bank construction can be decomposed into two main stages: Question Characteristics Analysis and Question Bank Development: (1) Question Characteristics Analysis first detailedly examines the properties and attributes of potential questions. Then, (2) Question Bank Development assembles the final question bank $\mathcal{Q}$ from the analyzed questions.
6.1 Question Characteristics Analysis
@@
The first stage, question characteristics analysis, involves a detailed examination of the properties and attributes of potential questions, e.g., difficulty, discrimination, and the knowledge concepts required to answer the question. For example, when selecting questions based on Fisher Information, one must leverage pre-calibrated parameters like difficulty ($\beta_{j}$), discrimination ($\alpha_{j}$), and guessing factor ($c_{j}$), alongside the current proficiency estimate, to compute the Information value $\mathcal{I}_{j}(\theta)$ for each question $j$. The methods of characteristics analysis can be categorized into three main approaches: expert-based, statistic-based, and deep learning-based methods.
Expert-based Characteristics Annotation. In expert-based annotation, domain experts assess question parameters, as seen in online CAT systems like SIETTE [36] and GenTAI [104]. Effective expert estimation often involves structured questionnaires [105], followed by discussions to resolve divergent opinions. Results are aggregated using averages for continuous attributes or voting for discrete ones [84]. Expert judgments can be subjective, leading to potential inaccuracies, especially with limited or inconsistent expert input. With the advancement of generative AI, LLMs can also be used to annotate question characteristics [103].
@@
Statistic-based Characteristics Annotation. The statistic-based method for annotating question characteristics requires gathering responses from a large group of examinees. It is resource-intensive nature and involves pre-testing with examinees [39]. In Classic Test Theory, question difficulty is calculated as the proportion of correct responses within examinees [109, 43], while discrimination is derived from performance disparities between higher and lower ability examinees [26]. The Q-matrix is another crucial characteristic of questions. It is a binary matrix that indicates which knowledge concepts are associated with a question. Numerous researchers have attempted to employ some parameter estimation approaches (e.g., maximum likelihood estimation and Bayesian estimation), to learn these characteristic parameters from response data [101, 147, 175].
@@
Deep Learning-based Characteristics Annotation. With the rise of Natural Language Processing (NLP), there has been an increasing trend in recent years to directly use the textual information of questions to analyze various attributes. For difficulty prediction, attention-based CNN models and domain adaptation strategies have been used to evaluate reading questions and medical question complexity [77, 126, 75]. For knowledge concept (Q-matrix) prediction, which typically exhibits a hierarchical structure, a Hierarchical attention-based Recurrent Neural Network has been proposed [73, 74]. Lei et al. [88] further take into account the multi-modal features of questions, such as images and formulas. Pre-trained NLP models have also proven effective for automated question analysis [178, 120].
6.2 Question Bank Development
@@
The second stage, question bank development, involves the actual assembly of the question bank from the analyzed questions of the first stage. This process should aim to create a balanced and varied bank that can cater to different levels of proficiency and different areas of knowledge [129, 128, 142]. According to different scenarios, the approaches to developing a question bank can be categorized into the following three aspects.
@@
Question Bank Blueprint Design. The goal of the blueprint design is to create an optimal framework for a bank, outlining the distribution of questions based on various attributes. Reckase et al. [128] analyze the characteristics of an optimal question bank in a CAT system using a 1PL-IRT model with a maximum Fisher information selection algorithm. They propose the bin-and-union method to allow a maximum deviation $r$ between optimal difficulty and estimated proficiency, extending these methods for large-scale CAT systems and continuous new question pretesting [66, 155].
@@
Question Bank Assembly. While the blueprint design focuses on creating an optimal framework, the assembly process involves generating question banks from an existing master bank according to specific requirements. Way et al. [173] discussed the development and maintenance of a master bank, including constraints to ensure the assembled question bank meets desired specifications. A mixed-integer programming [153] was proposed to create a bank that satisfies content specifications and maximizes information at selected proficiency values.
@@
Question Bank Rotating. Rotating the question bank involves dividing a master bank into smaller banks with overlapping elements, ensuring balanced exposure rates [146, 5, 149]. Ariel et al. [5] proposed dividing a master bank into smaller banks using Gulliksen’s matched random subtests method [64] to prevent over- or underexposure. The Weighted Deviation Model [149] manages the degree of overlap, maintaining representativeness and preventing question overexposure.
@@
Discussion: Think of the entire bank development process as creating and managing a library. The blueprint design is like the architectural plan for the library, defining where each section (e.g., fiction, non-fiction) will be located; The assembly process resembles acquiring books from suppliers based on specific demands; Rotating the question bank is similar to periodically rotating the books on display. Even though the library has a vast collection, only a subset is displayed prominently at any given time. This rotation ensures that different books get exposure, and library visitors encounter a variety of books over time. Although this section has so far focused on classical methods, the bank construction pipeline can also incorporate LLMs as auxiliary components. In other words, LLMs can be integrated into the bank development stage to improve scalability and reduce manual cost, while the psychometric principles of CAT remain unchanged. In this analogy, LLMs (or agents) can be viewed as “librarians” that help scale and accelerate curation: they can draft candidate items on demand under explicit constraints, and generate useful metadata (topic tags, expected solution outlines, and common error patterns) that supports indexing and retrieval. The construction of a high-quality question bank introduced in this section is not just a prerequisite for CAT, but also a continuous process. It requires regular updates and refinements to ensure the relevance and effectiveness of the adaptive testing system.
7 Test Control of CAT